OpenStudy (anonymous):

A population doubles every 21 years. Assuming exponential growth find the following: (a) The annual growth rate (b) The continuous growth rate per year Plz help, very many thanks math gods

OpenStudy (paxpolaris):

what's the formula for exponential growth?

OpenStudy (paxpolaris):

\[\Large P=P_0(1+r)^t\] population, \(P\) is given by this formula , where \(P_0\) is initial \(r\) is the annual growth rate , \(t\) is time in years

OpenStudy (paxpolaris):

since it doubles every 21 years:\[\large 2\cancel{P_0}=\cancel{P_0}(1+r)^{21}\]

OpenStudy (paxpolaris):

\[\ln2=21\ln(1+r)\] r=\[r\\=e^{\ln2/21}-1\\=2^{1/21}-1\\ \approx3.356\ \%\]

OpenStudy (perl):